Getting reaquainted with the OA system
Oysters are set to go into pH treatments (ambient, 7.5) on Wedneday 2/8, so on Sunday I went to Manchester to get the OA system running, so we could test the CO2 injectors and Durafets.
- Steven had opened valves on Saturday, so when I arrived I confirmed that pumps worked, float valves worked, and temperature was 10.3 & 10.4 in the headers, and ~10.8 in the culture tank. This, according to the Durafet’s temperature sensors.
- Untangled Durafets, and installed one in each header tank, and one in a culture tank.
- “Pink” in Header 1
- “Yellow” in Header 2
- “Green” in culture tank connected to Header 1
“Blue” Durafet was malfunctioning; error on control panel showed code “error input pv 1”. Did some user manual research, and a good cleaning might be necessary. Also, Rhonda provided very helpful info on GitHub Issue 467
- Opened the CO2 valve, and adjusted the regulator to provide for enough/not too much back pressure; setting is currently 50 psi
- Programmed both Header 1 & Header 2 to maintain pH setpoints of 7.45-7.55; injections are 0.8 seconds in duration, every 100 seconds (if needed)
- Plugged in algae dosing pumps to see if they work - they do.
- Identified appropriate tubing for the dosing pumps: 0.250”ID x 0.375”OD x 0.062” Wall Kuri Tec® Series 220 Low Density Food Grade Tubing, Polyethylene, Natural, 500’ Coil Item # 220-0662X500 from Harrington
To Do on Wednesday:
- Install new valves leading to culture tanks
- Insulate last 3 tanks
- Trouble shoot faulty Durafet, as per Rhonda’s suggestions
- Check continuous data from Durafets in both header tanks to see if pH 7.5 has been maintained. Determine how to adjust settings, if needed.
- Decide which header tank to use first for CO2 treatment
- Drain headers, refill (need to do both, so temperatures are the same)
- Adjust settings on low pH header; it likely has been getting too much CO2
- Put animals in new tanks
- Install algae dosing pumps (if not already)
- Decide if we want to start them on pH treatments that day, or possibly wait until Saturday (if we don’t get the continuous data from Rick, may want to wait…)
- Review chemistry sampling protocol; Yaamini needs to meet with Safety Dave if she is using HgCl